University of Wisconsin-Madison

South Pole Weekly Report: Week ending December 26


Population and Deployments

South Pole Station Population is currently at 230 and there are 38 IceCube staff on the ice. Five IceCubers took off last week, and three new arrivals during the week including the new on-ice lead. The five departures for the week unfortunately got stuck at McMurdo Station over the holiday weekend.

Those of us here at Pole celebrated Christmas in the usual way. The day started with the traditional Race Around the World. A course is laid out around the South Pole station and participants "race" around the world setting foot in all 24 worldwide time zones. We then had a delicious Christmas dinner with main courses of beef wellington and lobster tails. All Cubers on station then gathered in the evening and waited for Santa to appear. It seems the IceCube team must have been good since everyone received something special. It was a great day for us though we dearly miss our families and friends back home.

Cargo is on schedule for IceCube. We will begin moving cargo north this week with our first scheduled shipment on Tuesday.

Drilling and Deployment

With drilling over, we have moved into the storage phase of the season. The Enhanced Hot Water Drill is being prepped for storage, and this week we will drain the water from the seasonal equipment site and begin removing hoses and plumbing. Crews are working on removing the drill hose from the main hose reel and spooling it up for storage at Pole, work is ongoing to prep the drill towers for storage. Work is continuing to finish up some improvements on the EHWD.

There are no more string deployments to report. An inventory was made of remaining string deployment materials, which could be of potential use in a future detector installation. Those materials have been put in a large wooden crate and will be stored together with the hot water drill. They include things like several hundred shackles and around 20 of the large Yale grips. A significant number of steel weights (1 to 2 tons) used for the strings will also stay here.


Special devices

The radio pulser and Dark Matter Ice (DM-Ice) units that were co-deployed with IceCube strings await full- freeze in for further testing. The Swedish Camera continues to acquire very interesting images as hole 80 freezes in, such as the images seen here, showing the spacer rings between the two cameras are partially engulfed by ice, eight days after deployment.

The Paro sensor logging system continues to monitor conditions at the bottom of holes 80, 22, 14 and 7.

IceTop

We are now in the monitoring phase of the freeze process. It is important to keep the drifting snow off of the ice surface or it will act as an insulator and slow the freezing. We have had a lot of high winds lately that caused drifting in the tanks. Thanks to those who did an excellent job clearing snow from the ice surface. All of the DOM hanger boards were removed to make snow removal easier; the boards have been disassembled and will be scrapped.


Monitoring shows that dissolved gasses in the IceTop tanks is minimal so the duty cycle of the circulation pumps has been backed down to 25% in most tanks. This reduces the heat added to the water by the pump and will help speed freezing. We made modifications to the data transfer script to improve the reliability of getting our data north.

Inside the IceCube Lab

This week we completed the new web sever which contains our mediawiki, printing software and the winter-over Wiki. All new Dell servers installed in the IceCube Lab are now in the IT monitor software. We have also successfully completed all the data aquisition machines. We have started the cleanup and retro phase, we are going to start packing up old and failed equipment from the past winter and start sending it back north, this will also include cleaning and organizing the stockroom in the lab.

Other news


It was a beautiful Christmas Day, sunny and with no wind--perfect conditions for the Race Around the World.